A 1.0 kg ball and a 4.0 kg ball are connected by a 3.0-m-long rigid, massless rod. The rod and balls are rotating clockwise about its center of gravity at 25 rpm. 


Part A 

What magnitude torque will bring the balls to a halt in 5.0 s?  
Answer #1

Center of gravity :

X = 4*3/5

X = 2.4 m

Moment of inertia :

I = 1*2.4^2 + 4*0.6^2

I = 7.2 kg m^2

Angular velocity :

w = 25*2*3.14/60

w = 2.61 rad/s

Angular acceleration :

a = 0 - 2.61/5

a = - 0.523 rad/s^2

Torque :

T = I*a

T = 7.2*(-0.523)

T = - 3.768 Nm